# Quadruped robot traversing 3D complex environments with limited perception

### v1

@@ I-A Related Works
Existing collision detection methods can be primarily classified into model-based methods or model-free methods. [2], [3], [4], [5] detect collision by comparing the estimated torques with predefined threshold. Among these, [2] proposes a momentum-based disturbance observer method. In comparison to the GM method, the MESO method introduced by [2] demonstrates improved collision force estimation under similar noise levels in practical systems. Both methods can locate the robot link with collision and provide directional information on the Cartesian collision force. However, in practical robot applications, these model-based methods require high precision in the robot’s model and are sensitive to disturbances and loads.
@@
Recently, methods for collision detection, isolation, and identification in humanoid robots have been introduced [11, 12, 13]. In [13], relying solely on proprioceptive sensing, the authors combine and compare the GM approach with an F/T-sensor-based approach. However, these methodologies have predominantly been tested in simulated environments with flat terrain. In real-world applications with robots, additional challenges such as communication delays, sensor synchronization, and noise must be taken into account. In the realm of body collision detection for quadruped robots, prior model-based efforts have proposed various state estimation techniques for the robot[14], [15], [16], attempting to fuse different sensors for more accurate state estimation [17]. Under conditions where only onboard perception is available, Fink et al. propose a sensor-less model based on kinematics to estimate the location of a single contact point at the shin level in real robot[10].
@@ II-B Base Set
We model the environment as a Markov Decision Process (MDP). An MDP is defined by a tuple$(S,A,P_{a},R_{a})$, where $S$ represents the set of all possible states, and $A$ represents the action space, $P(s_{t+1}|s_{t},a_{t})$ is the state transition function, indicating the probability of transitioning to state $s_{t+1}$ after taking action $a_{t}$ in state $s_{t}$, and $R(s_{t+1}|s_{t},a_{t})$ is the immediate reward. The agent selects an action $a_{t}$ from the policy based on the current state $s_{t}$. For the state $s_{t}$ and action $a_{t}$, the MDP calculates the next state $s_{t+1}$ and reward $r_{t}$, and then provides feedback to the agent. The goal is to select a policy that maximizes the cumulative function of random rewards, expressed as:
@@
State Space: For teacher policy, the state space $s_{t}$ is composed of proprioceptive observation $o_{t}$, explicit observations $\hat{c}_{t}$, $\hat{v}_{t}$, where $\hat{c}_{t}$ is the estimate of collision information, $\hat{v}_{t}$ is the estimate of body linear velocity, and implicit observations $p_{t}$, $e_{t}$, refer to the latent variables obtained after the 3D collision domain and privileged information are processed through the encoders. Privileged observation $e_{t}$, including body mass, center of mass,friction and motor strength. the state space $s_{t}$ as shown in the following equation:
@@
The structure of student policy is kept consistent with the teacher policy, where $e_{t}$ and $p_{t}$ are estimated using proprioceptive observations. Specifically, in the first phase, we employ Regularized Online Adaptation (ROA)[42] to train a privileged information estimation module. In the second phase, we obtain estimates of privileged information $\hat{e}_{t}$. Furthermore, in the second phase, we also train a hybrid imagination module to obtain estimates of $\hat{p}_{t}$, as detailed in section II.D.
@@
The linear velocity reward with heading constraint ensures that the robot moves along the direction of the obstacle placement, thus preventing the robot from choosing to bypass the obstacle. This design of $L_{d_{v}}$ stems from observations made during the training process: when faced with difficult obstacles, robots tend to adopt a policy of quickly bouncing back to their original position, then approaching the obstacle at a set speed and bouncing back again, which easily leads to the policy falling into a local optimum. Therefore, we included a penalty for negative velocity in the speed tracking reward to prevent the robot from adopting a rebound policy.Where both $k_{positive}$ and ${k_{negative}}$ are constant quantities, positive reward coefficient and negative penalty coefficient, respectively,
To enable the robot to sidestep through narrow spaces and maintain a natural walking posture in normal environments, we introduced the following reward mechanism:
@@
The pos reward consists of two parts, the GuidePos reward does not directly limit the angle of the robot’s hip joint but penalizes the sum of the angles of the left and right leg hips to be zero and the difference in angles between the front and back leg hips, ensuring that the robot’s limbs can always walk perpendicular to the ground under any circumstances. This not only facilitates the generation of sidestepping actions but the NaturalPos reward ensures that the robot maintains a natural walking posture in normal environments. $w_{1},w_{2}$ corresponds to the weights of the two-part rewards.
@@ II-D Implicit Collision Domain imagination
To better perceive and respond to obstacles in the complex 3D environments, this study introduces the concept of a collision domain, which is an area around the robot’s body containing information about obstacles, as shown in Fig. 3. Specifically, we have set a cuboid sampling domain which is centered on the robot, see Table 1 for specific values, within which the distance between the robot and obstacles is sampled to detect imminent collisions. In complex environments such as dense forests and low visibility areas where external sensors like lidar and cameras often fail, proprioception becomes the sole means of understanding the collision domain. Therefore, this study proposes a method that uses only proprioception to deeply estimate the characteristics of the collision domain.
@@ III-D Simulation Experiments
Baselines and Ablations. We conducted a comparative analysis of our method against baseline approaches such as RMA and WTW, which are two typical methods in this domain. The RMA approach extracts information from 2D elevation maps and privileged data to adapt to the environment, whereas WTW relies on expert knowledge to set rewards and manually adjust the robot’s posture for obstacle traversal. In addition to RMA’s framework, we implemented our reward function. As shown in TABLE III, the 2D elevation maps used in RMA struggle to capture the complete features of 3D obstacles in the environment, resulting in significantly lower performance across multiple tasks compared to our method. WTW can control the robot’s height, which allows the robot to crawl to traverse Tunnel obstacles but lacks the ability to climb or sidestep, failing to navigate complex obstacles like Highland, Crack, and Barrier.
@@ III-E Real-World Experiments
Quantitative Experiment. As depicted in the Fig. 8., we have quantitatively compared our method with three other strategies: Go2 Default, Go2 Special, and WTW, in real-world scenarios. The data shown represents the average success rate of the robot when confronting various types of obstacles at different levels of difficulty, across 10 trials, aiming to showcase the obstacle negotiation capability of our approach in actual environments. It is clear from the figure that experiments utilizing our policy exhibited superior performance in almost all obstacle types. Notably, in the specific scenario of navigating through Tunnels, the WTW policy performed slightly better, primarily due to its use of human knowledge priors and manual adjustment of the robot’s height for successful passage. In contrast, my method is adaptive, which may lead to certain errors in estimating environmental obstacles, resulting in a degree of performance degradation. However, this performance degradation is within an acceptable range.
@@ IV-A Reward Function
Reward items are listed in TABLE III. The task reward helps the robot adapt to the specific task scenarios, and the details of the design (4,6,7) is shown in II.B above. Where we set the Goal Velocity’s reward coefficient $k_{positive}$ to 1 and the penalty coefficient $k_{negative}$ to -3. This set of parameters was optimized based on our comprehensive simulation experiments. Additionally, we designed the regularization reward referring to [46], adjusting it according to the type of robots and task scenarios we are using.
@@ IV-B Additional Training Details
Hybrid Imagination Model $\boldsymbol{o}_{t-10},\hat{\boldsymbol{c}}_{t-10}...,\hat{\boldsymbol{c}}_{t},\boldsymbol{o}_{t}$ / $\hat{\boldsymbol{p}}_{t}$
@@
Policy training and Imitation training We use PPO with hyperparameters listed in TABLE V to train the teacher policy. We regard the process of teacher policy supervising student policy learning as imitation learning process, $a_{\text{teacher}}$ and $a_{\text{student}}$ are the action vectors from the teacher and student respectively in an actor network. The gradient of the loss $\mathcal{L}$ can be defined as the sum of the squared norms of the difference between $a_{\text{teacher}}$ and $a_{\text{student}}$ over all actions:

### v3

@@ I-A Related Works
Existing collision detection methods can be primarily classified into model-based methods or model-free methods. [2], [3], [4], [5] detect collision by comparing the estimated torques with predefined threshold. Among these, [2] proposes a momentum-based disturbance observer method. In comparison to the GM(generalized momenta-based) method, the MESO (modified extended state observer) method introduced by [5] demonstrates improved collision force estimation under similar noise levels in practical systems. Both methods can locate the robot link with collision and provide directional information on the Cartesian collision force. However, in practical robot applications, these model-based methods require high precision in the robot’s model and are sensitive to disturbances and loads.
@@
Recently, methods for collision detection, isolation, and identification in humanoid robots have been introduced [11, 12, 13]. In [13], relying solely on proprioceptive sensing, the authors combine and compare the GM approach with an Force/Torque-sensor-based approach. However, these methodologies have predominantly been tested in simulated environments with flat terrain. In real-world applications with robots, additional challenges such as communication delays, sensor synchronization, and noise must be taken into account. In the realm of body collision detection for quadruped robots, prior model-based efforts have proposed various state estimation techniques for the robot[14], [15], [16], attempting to fuse different sensors for more accurate state estimation [17]. Under conditions where only onboard perception is available, Fink et al. propose a sensor-less model based on kinematics to estimate the location of a single contact point at the shin level in real robot[10].
@@ II-B Base Set
We model the environment as a Markov Decision Process (MDP). An MDP is defined by a tuple$(S,A,P_{a},R_{a})$, where $S$ represents the set of all possible states, and $A$ represents the action space, $P(s_{t+1}|s_{t},a_{t})$ is the state transition function, indicating the probability of transitioning to state $s_{t+1}$ after taking action $a_{t}$ in state $s_{t}$, and $R(s_{t+1}|s_{t},a_{t})$ is the immediate reward. The agent selects an action $a_{t}$ from the policy based on the current state $s_{t}$. For the state $s_{t}$ and action $a_{t}$, the MDP calculates the next state $s_{t+1}$ and reward $r_{t}$, and then provides feedback to the agent. The goal is to select a policy that maximizes the cumulative sum of discounted rewards, expressed as:
@@
State Space: For teacher policy, the state space $s_{t}$ is composed of proprioceptive observation $o_{t}$, explicit observations $\hat{c}_{t}$, $\hat{v}_{t}$, where $\hat{c}_{t}$ is the estimate of collision information, $\hat{v}_{t}$ is the estimate of body linear velocity, and implicit observations $p_{t}$, $e_{t}$, refer to the latent variables obtained after the 3D collision domain and privileged information are processed through the encoders. Privileged implicit observation $e_{t}$ includes encoded body mass, center of mass, friction and motor strength. The state space $s_{t}$ is shown in the following equation:
@@
The structure of student policy is kept consistent with the teacher policy, where $e_{t}$ and $p_{t}$ are estimated using proprioceptive observations. Specifically, in the first phase, we employ Regularized Online Adaptation (ROA)[42] to train a privileged information estimation module. In the second phase, we obtain estimates of privileged implicit information $\hat{e}_{t}$. Furthermore, in the second phase, we also train a hybrid imagination module to obtain estimates of $\hat{p}_{t}$, as detailed in section II.D.
@@
The linear velocity reward with heading constraint ensures that the robot moves along the direction of the obstacle placement, thus preventing the robot from choosing to bypass the obstacle. This design of $L_{d_{v}}$ stems from observations made during the training process: when faced with difficult obstacles, robots tend to adopt a policy of quickly bouncing back to their original position, then approaching the obstacle at a set speed and bouncing back again, which easily leads to the policy falling into a local optimum. Therefore, we included a penalty for negative velocity in the speed tracking reward to prevent the robot from adopting a rebound policy.Where both $k_{positive}$ and ${k_{negative}}$ are constant quantities, positive reward coefficient and negative penalty coefficient respectively, the detailed values are in the appendix. To enable the robot to sidestep through narrow spaces and maintain a natural walking posture in normal environments, we introduced the following reward mechanism:
@@
The pos reward consists of two parts, the GuidePos reward does not directly limit the angle of the robot’s hip joint but penalizes the sum of the angles of the left and right leg hips to be zero and the difference in angles between the front and back leg hips, ensuring that the robot’s limbs can always walk perpendicular to the ground under any circumstances. This not only facilitates the generation of sidestepping actions but the NaturalPos reward ensures that the robot maintains a natural walking posture in normal environments. $W_{1},W_{2}$ corresponds to the weights of the two-part rewards, the detailed values are in the appendix.
@@ II-D Implicit Collision Domain imagination
To better perceive and respond to obstacles in the complex 3D environments, this study introduces the concept of a collision domain, which is an area around the robot’s body containing information about obstacles, as shown in Fig. 3. Specifically, we have set a cuboid sampling domain centered on the robot, see Table 1 for specific values. A boolean value indicates whether each point of the grid is inside an obstacle to detect imminent collisions. In complex environments such as dense forests and low visibility areas where external sensors like lidar and cameras often fail, proprioception becomes the sole means of understanding the collision domain. Therefore, this study proposes a method that uses only proprioception to deeply estimate the characteristics of the collision domain.
@@ III-D Simulation Experiments
Baselines and Ablations. We conducted a comparative analysis of our method against baseline approaches such as RMA and WTW, which are two typical methods in this domain. The RMA approach extracts information from 2D elevation maps and privileged data to adapt to the environment, whereas WTW relies on expert knowledge to set rewards and manually adjust the robot’s posture for obstacle traversal. In addition to RMA’s framework, we implemented our reward function. As shown in TABLE II, the 2D elevation maps used in RMA struggle to capture the complete features of 3D obstacles in the environment, resulting in significantly lower performance across multiple tasks compared to our method. WTW can control the robot’s height, which allows the robot to crawl to traverse Tunnel obstacles but lacks the ability to climb or sidestep, failing to navigate complex obstacles like Highland, Crack, and Barrier.
@@ III-E Real-World Experiments
Quantitative Experiment. As depicted in the Fig. 8., we have quantitatively compared our method with three other strategies: Go2 Default, Go2 Special, and WTW, in real-world scenarios. The data shown represents the average success rate of the robot when confronting various types of obstacles at different levels of difficulty, across 10 trials, aiming to showcase the obstacle negotiation capability of our approach in actual environments. It is clear from the figure that experiments utilizing our policy exhibited superior performance in almost all obstacle types. Notably, in the specific scenario of navigating through Tunnels, the WTW policy performed slightly better, primarily due to its use of human knowledge priors and manual adjustment of the robot’s height for successful passage. In contrast, our method is adaptive, which may lead to certain errors in estimating environmental obstacles, resulting in a degree of performance degradation. However, this performance degradation is within an acceptable range.
@@ IV-A Reward Function
Reward items are listed in TABLE III. The task reward helps the robot adapt to the specific task scenarios, and the details of the design (4,6,7) is shown in II.B above. Where we set the Goal Velocity’s reward coefficient $K_{positive}$ to 1 and the penalty coefficient $K_{negative}$ to -3. This set of parameters was optimized based on our comprehensive simulation experiments. Additionally, we designed the regularization reward referring to [46], adjusted it according to the type of robots and task scenarios.
@@ IV-B Additional Training Details
Hybrid Imagination Model $\hat{\boldsymbol{c}}_{t},\boldsymbol{o}_{t}$ [128, 64] $\hat{\boldsymbol{p}}_{t}$
@@
Policy training and Imitation training We use PPO with hyperparameters listed in TABLE V to train the teacher policy. We regard the process of teacher policy supervising student policy learning as imitation learning process, $a_{\text{t}}$ and $\hat{a}_{\text{t}}$ are the action vectors from the teacher and student respectively in an actor network. The gradient of the loss $\mathcal{L}$ can be defined as the sum of the squared norms of the difference between $a_{\text{t}}$ and $\hat{a}_{\text{t}}$ over all actions: